An electron is moving east at 5x106 m/s. Determine the magnetic field from a point north east with a distance of 5cm. 



1
Expert's answer
2021-05-07T09:47:24-0400

"\bar{B}= \frac{\mu_0}{4 \pi} \frac{qvsin \theta}{r^3}"

"\bar{B}= \frac{1.2566 *10^{-6}}{4 \pi} \frac{1.60217662 *10^{-19}*5*10^6*sin45}{(0.05)^3}"

"\bar{B}=5.45303 \times 10^{-16} T"